# Data-driven Forecasting of Deep Learning Performance on GPUs

### v1

@@ 6.2. Results
End-to-end mean percentage error. Figure 7 shows the inference and training latency prediction error of NeuSight and the baselines on 6 deep learning models with varying batch sizes executed on 8 GPUs in percentage error. Instances that include out-of-distribution GPUs or models are highlighted. For example, H100 is an out-of-distribution GPU as it is not included in the training set. Similarly, GPT3-2.7B is an out-of-distribution model as it includes BMM with one of the operand dimensions equal to 2048, which is larger than the 1024 used in the training set.
@@
Across all GPUs and models, NeuSight shows a modest average percentage error compared to all other baselines. Specifically, the roofline analysis shows an average error of 32.4% and 31.8%, Habitat shows 79% and 259%, and Li et al. shows 62% and 59%, for inference and training, respectively. In contrast, NeuSight shows a percentage error of only 9.2% for inference and 8.1% for training workloads.
@@
The advantage of NeuSight over other baselines, especially Habitat and Li et al., becomes more apparent when evaluated on out-of-distribution GPUs. NeuSight can predict the latency of the unseen GPUs, demonstrating an average error of 8.7% and maximum error of 26.24%. Habitat and Li et al. struggle with out-of-distribution GPUs, show an average error of 232.8% and 88.7%, and maximum error of 2067.7% and 425.3% respectively, even though they are trained on the same data as NeuSight. The main reason for the lower absolute percentage error of NeuSight is that it does not directly use the MLP to make latency predictions, instead applies GPU architecture and execution insights atop the machine learning predictions. In the case of newer generations GPUs like H100, which exhibit features outside the distribution seen in the training set, Habitat’s MLP prediction encounters challenges in accurately predicting GEMM operators in DNN workloads. Li et al. also falls short of NeuSight due to its limited method of inferring hardware utilization of unseen GPUs. This superior performance highlights the robustness of NeuSight in handling out-of-distribution features and making accurate predictions even on unseen GPUs.
@@
BERT-Large 8 Measured Latency (ms) 409.5 344.1 212.1 195.0 69.5 65.0
NeuSight Prediction (ms) 338.4 (17.4%) 264.3 (23.2%) 213.9 (0.8%) 198.3 (1.7%) 87.8 (26.2%) 80.7 (24.3%)
16 Measured Latency (ms) 877.1 724.7 408.5 385.2 136.6 127.0
NeuSight Prediction (ms) 645.6 (26.4%) 519.5 (28.3%) 405.6 (0.7%) 378.2 (1.8%) 170.3 (24.7%) 157.6 (24.1%)
GPT2-Large 4 Measured Latency (ms) 1242.2 886.8 550.1 484.0 213.6 174.7
NeuSight Prediction (ms) 1080.0 (13.1%) 849.0 (4.3%) 571.3 (3.9%) 525.5 (8.6%) 208.0 (2.6%) 186.8 (6.9%)
8 Measured Latency (ms) 2485.4 1775.1 1110.1 959.4 411.9 343.4
NeuSight Prediction (ms) 2278.0 (8.3%) 1816.9 (2.4%) 1169.2 (5.3%) 1083.6 (12.9%) 468.4 (13.7%) 429.0 (24.9%)
@@
Batch A100-40GB x 4 (NVLink) H100 x 4 (DGX Box)
Size Data Parallel Tensor Parallel Pipeline Parallel Data Parallel Tennsor Parallel Pipeline Parallel
GPT2-Large 8 Measured Latency (ms) 466.4 460.2 807.4 202.1 206.4 341.1
NeuSight Prediction (ms) 461.2 (1.1%) 461.0 (0.2%) 784.8 (2.8%) 195.0 (3.5%) 198.0 (4.1%) 330.7 (3.0%)
16 Measured Latency (ms) OOM OOM OOM 627.8 688 1146.6
NeuSight Prediction (ms) 603.9 (3.8%) 805.3 (17.0%) 1018.5 (11.2%)
GPT3-XL 4 Measured Latency (ms) OOM OOM OOM OOM 987.8 1715.7
NeuSight Prediction (ms) 1001.7 (1.4%) 1521.9 (11.3%)
Per-operator prediction error. Figure 8 shows the prediction error for each kernel type (BMM, fully-connected, element-wise, softmax, layer normalization), averaged across all workloads evaluated in Figure 7. The roofline analysis consistently shows a higher percentage error compared to NeuSight for both in-distribution and out-of-distribution data points, with an average percentage error of 34%. For, Habitat and Li et al. their accuracy significantly degrades for out-of-distribution data points, especially for BMM and fully-connected operators. Habitat and Li et al. show percentage error of 135.9% and 29.6% for BMM, and 264.7% and 148.2% for fully-connected operators, respectively. In contrast, NeuSight shows better accuracy than roofline analysis, Habitat and Li et al., with a percentage error of 11.9% for BMM operators and 15.8% for fully-connected operators.
Operator fusion. Table 5 shows the inference latency prediction error of BERT-Large and GPT2-Large with operator fusion on L4, A100 (40GB), and H100. Using torch.compile (PyTorch, b), the built-in operator fusion tool in PyTorch, we fused the supported operators and measured the latency of the fused models. NeuSight predicts latency with a 13.6% error across all fused models, 17.2% for BERT-Large, and 10% for GPT2-Large.
@@
Distributed Execution. Table 6 shows the training latency predictions for distributed execution on two types of GPU servers: one with 4 A100s-40GB connected via NVLink and 4 H100s in a DGX Box (NVIDIA, 2023b), training GPT2-Large and GPT3-XL. The A100 GPUs are connected in a mesh topology with each GPU using 12 NVLinks, providing a bi-directional bandwidth of 600GB/s. In the H100 DGX system, each GPU is connected with 18 NVLinks, providing a bi-directional bandwidth of 900GB/s. Both configurations enable full-bandwidth communication between any two GPUs. We evaluate NeuSight on data, tensor model, and pipeline parallel execution, albeit individually. This is because NeuSight aims to also estimate the latency for a wide range of models, including the ones which cannot fit on a single device due to memory, thus require tensor model or pipeline to split the DNN graph. For each method, we split the execution across the 4 GPUs. For pipeline parallelism, we have 4 micro-batches. Across all the models, NeuSight predicts the latency of distributed training with an average error of 5.4%, including 6.9% for the H100 server and 1.4% for the A100 server. This demonstrates that NeuSight predicts effectively for per-kernel, per-GPU, and distributed execution on various GPUs and server types. This is mainly due to the robust underlying mechanism of kernel estimation, that can be extended with the network level estimates to forecast distributed execution latencies.
@@ 7. Conclusion
In this work, we present a novel framework, NeuSight, that forecasts the performance of deep learning execution on GPU servers. In developing NeuSight, we leverage the execution deep learning kernel on a GPU architecture such as tiled execution and use this insight to decompose the prediction problem to smaller sub-problems, i.e., per-tile prediction. For each tile, we use machine learning models to capture the complex non-linear relationship between achieved throughput and input features, and impose performance laws to bound the predicted latency. Results on a variety of models and GPUs, show that NeuSight can provide better predictions than prior works in the area. We will open-source NeuSight to benefit broader research community.

### v2

@@ 6.2. Single-Device Execution Results
End-to-end mean percentage error. Figure 7 shows the inference and training latency prediction error of NeuSight and the baselines on 6 deep learning models with varying batch sizes executed on 8 GPUs in percentage error. Instances that include out-of-distribution GPUs or models are highlighted. For example, H100 is an out-of-distribution GPU as is not included in the training set. Similarly, GPT3-2.7B is an out-of-distribution model as it includes BMM with one of the operand dimensions equal to 2048, which is larger than the 1024 used in the training set.
@@
Across all GPUs and models, NeuSight shows a modest average percentage error compared to all other baselines. Specifically, the roofline analysis shows an average error of 31.2% and 31.9%, Habitat shows 220.9% and 725.8%, and Li et al. shows 61.2% and 58.3%, for inference and training, respectively. In contrast, NeuSight shows a percentage error of only 9.7% for inference and 7.3% for training workloads. Smaller models relative to GPU capacity, like BERT on H100, tend to show higher error rates due to library overheads. Nonetheless, NeuSight’s kernel tiling approach still achieves reasonable accuracy.
@@
The strengths of NeuSight are highlighted on out-of-distribution GPUs, where it achieves an average latency prediction error of 8.1% and a maximum error of 28.2%. In comparison, Habitat and Li et al. show average errors of 724.3% and 94.0%, with maximum errors of 4529.9% and 435.9%, respectively, when trained on the same dataset with NeuSight. The main reason for the lower absolute percentage error of NeuSight is that it does not directly use the MLP to make latency predictions, instead applies GPU architecture and execution insights atop the machine learning predictions. In the case of newer generations GPUs like H100, which exhibit features outside the distribution seen in the training set, Habitat’s and Li et al. observe higher errors. The lower error rate of NeuSight showcases its robustness in handling out-of-distribution features and making accurate predictions even on unseen GPUs.
@@
Per-operator prediction error. Figure 8 shows the prediction error for each kernel type (BMM, fully-connected, element-wise, softmax, layer normalization), averaged across all workloads evaluated in Figure 7. The roofline analysis consistently shows a higher percentage error compared to NeuSight for both in-distribution and out-of-distribution data points, with an average percentage error of 34%. The accuracy of Habitat and Li et al. degrades on out-of-distribution data points, particularly for BMM and fully-connected operators. Habitat and Li et al. show percentage error of 123.2% and 30.0% for BMM, and 799.3% and 152.6% for fully-connected operators, respectively. In contrast, NeuSight shows better accuracy than roofline analysis, Habitat and Li et al., with a percentage error of 13.8% for BMM operators and 13.9% for fully-connected operators.
Table 6 shows the contribution of each layer type to the end-to-end latency. Since GEMM operations contribute the most to the overall latency, accurate GEMM predictions of NeuSight are crucial for maintaining low error rates in end-to-end performance predictions. We observed higher error rates for layer normalization due to its generally short latency, which makes accurate predictions challenging. However, as layer normalization contributes only up-to 2% to the overall latency, its higher error does not significantly affect NeuSight’s end-to-end prediction accuracy.
Table 6. Per-operator contribution to the overall latency for various models on H100 for model inference. The contributions from different types of operators are shown as percentages.
Model Batch Size BMM LINEAR EW SOFTMAX LN OTHERS
BERT-Large 16 12% 74% 10% 3% 2% 0%
GPT2-Large 4 11.8% 62.8% 14.9% 4.0% 1.1% 5.3%
OPT13 2 13.1% 62.9% 9.6% 5.9% 0.6% 7.8%
GPT3-XL 2 9.5% 76.4% 7.9% 2.5% 0.4% 3.3%
GPU across vendors (AMD). Modern GPUs, including those from AMD, share many execution and deep learning concepts discussed in this paper. For instance, AMD’s CDNA architecture uses multiple identical compute units as its core building blocks, similar to NVIDIA’s Streaming Multiprocessors (SMs), and AMD’s HIP programming model employs workgroups similar to CUDA’s thread blocks(AMD, 2024). We trained NeuSight using data from MI100 and MI210 and evaluated its prediction accuracy on the MI250. We evaluated the prediction accuracy of NeuSight across 5 model architectures varying batch sizes, including GPT3-XL, GPT3-2.7B, and OPT-1.3B which contain out-of-distribution operations. Figure 9 presents the AMD GPU evaluation results. Across all tested models, NeuSight achieves an average percentage error of 8.8% for inference and 15.7% for training, demonstrating its ability to generalize across different GPU vendors.
(a) Inference latency prediction error
(b) Training latency prediction error
Figure 9. The inference and training latency prediction percentage error of NeuSight for various models across different batch sizes on AMD MI100 and MI210 GPUs.
@@
BERT-Large 8 Measured Latency (ms) 409.9 344.1 205.5 174.8 69.8 65.0
NeuSight Prediction (ms) 350.1 (14.6%) 294.3 (14.5%) 216.7 (5.4%) 204.8 (17.2%) 86.2 (23.4%) 80.9 (24.6%)
16 Measured Latency (ms) 855.4 724.7 396.5 345.2 136.8 127.0
NeuSight Prediction (ms) 681.1 (20.4%) 573.4 (20.9%) 414.1 (4.5%) 391.5 (13.4%) 154.8 (13.2%) 144.7 (14.0%)
GPT2-Large 4 Measured Latency (ms) 1276.3 886.8 535.8 477.6 215.0 174.7
NeuSight Prediction (ms) 1135.9 (11.0%) 952.5 (7.4%) 576.2 (7.5%) 537.6 (12.6%) 208.4 (3.1%) 190.6 (9.1%)
8 Measured Latency (ms) 2563.0 1775.1 1084.7 948.5 414.3 343.4
NeuSight Prediction (ms) 2299.0 (10.3%) 1943.8 (9.5%) 1179.4 (8.7%) 1107.9 (16.8%) 442.5 (6.8%) 409.9 (19.4%)
Operator fusion. Table 7 shows the inference latency prediction error of BERT-Large and GPT2-Large with operator fusion on L4, A100 (40GB), and H100. Using torch.compile (PyTorch, 2024a), the built-in operator fusion tool in PyTorch, we fused the supported operators and measured the latency of the fused models. NeuSight predicts latency with a 15.7% error across all fused models, 18.9% for BERT-Large, and 12.5% for GPT2-Large. This implies the generalizability of NeuSight to capture the common optimizations employed for deep learning kernels, as it distills the execution to tile and waves.
Figure 10. Performance of (NxN) x (NxN) matrix multiplication on H100 GPUs using FP16 numeric type and Tensor Cores across different batch sizes.
New numerical type and hardware unit. Even in the case of the introduction of new hardware units and software changes, NeuSight can be adapted to predict the performance of unseen hardware units. This is because core concepts such as multiple identical processing units, tiled execution, and fundamental performance equations remain relevant. We demonstrate this adaptability in this experiment, where NeuSight is tested on a new numerical type (FP16) and a new hardware unit (Tensor Core). Figure 10 presents the latency predictions by NeuSight for FP16 matrix-multiplication on H100 Tensor Cores, illustrating the versatility of NeuSight on new hardware. To adapt to changes in hardware and software features, we adjust the input features, accounting for the reduced memory requirement due to lower precision and the higher peak FLOPS to reflect the enhanced compute capacity of the Tensor Core. NeuSight predicts the latency of the batched matrix multiplication with an average error rate of 13%, showcasing its reliability even with these hardware and software changes.
@@
Global A100-40GB x 4 (NVLink) H100 x 4 (DGX Box)
Batch Size Data Parallel Tensor Parallel Pipeline Parallel Data Parallel Tensor Parallel Pipeline Parallel
GPT2-Large 4 Measured Latency (ms) 452.3 484.9 1597.6 194.9 211.1 637.5
NeuSight Prediction (ms) 510.5 (12.9%) 525.6 (8.4%) 1761.6 (10.3%) 214.1 (9.9%) 228.1 (8.0%) 650.2 (2.0%)
16 Measured Latency (ms) OOM OOM OOM 642.1 729.7 2384.0
NeuSight Prediction (ms) 649.6 (1.2%) 825.1 (13.1%) 2638.4 (10.7%)
GPT3-XL 4 Measured Latency (ms) OOM OOM OOM OOM 1010.5 3554.4
NeuSight Prediction (ms) 1047.6 (3.7%) 3717.0 (4.6%)
6.3. Distributed Execution Results
@@
Single-server execution. Table 8 shows the training latency predictions for distributed execution on two types of GPU servers: one with 4 A100s-40GB connected via NVLink and 4 H100s in a DGX Box (NVIDIA, 2023b), training GPT2-Large and GPT3-XL. The A100 GPUs are connected in a mesh topology with each GPU using 12 NVLinks, providing a bi-directional bandwidth of 600GB/s. In the H100 DGX system, each GPU is connected with 18 NVLinks, providing a bi-directional bandwidth of 900GB/s. Both configurations enable full-bandwidth communication between any two GPUs. We evaluate NeuSight on data, tensor model, and pipeline parallel execution, albeit individually. This is because NeuSight aims to also estimate the latency for a wide range of models, including the ones which cannot fit on a single device due to memory, thus require tensor model or pipeline to split the DNN graph. For each method, we split the execution across the 4 GPUs, and use a single micro-batch.
Across all the models, NeuSight predicts the latency of distributed training with an average error of 7.7%, including 6.7% for the H100 server and 10.5% for the A100 server. This demonstrates that NeuSight predicts effectively for per-kernel, per-GPU, and distributed execution on various GPUs and server types. This is mainly due to the robust underlying mechanism of kernel estimation, that can be extended with the network level estimates to forecast distributed execution latencies.
Table 9. NeuSight prediction for multi-node distributed execution of GPT-3 on H100 GPUs.
# Nodes 1 4 384 768 3840
NeuSight Prediction (ms) 1514.9 1836.7 12028.3 12135.5 12564.6
Multi-node distributed execution. Extending beyond single-server execution, NeuSight can be used with networking simulators to predict multi-node execution where GPUs communicate across different servers. While simulating multi-server networks is well handled by tools such as ASTRA-Sim (Rashidi et al., 2020) and ns-3 (nsnam, 2024), NeuSight’s predictions can be integrated with these simulators to forecast multi-node distributed training. To demonstrate this capability, we provide estimation results for executing GPT-3 across 1 to 3840 nodes for a single iteration. The experimental infrastructure consists of nodes equipped with 8 H100 GPUs each, interconnected via a multi-level fat-tree network using InfiniBand switches, with a link bandwidth of 100 Gbps. Nodes are connected hierarchically across levels consisting of 4, 384, 768, and up to 3840 nodes. Each DGX server provides GPU-to-GPU bandwidth of 900GB/s (NVIDIA, 2023b). For the distributed training configuration, we use an 8-wide Tensor Model parallel setup within each server, while we use data parallelism across the nodes. The per-node batch size is set to 8, and the global batch size scales as 8 times the number of nodes. NeuSight handles per-device predictions, while the analytical network performance model from the previous subsection estimates the communication latency for AllReduce operations across the switched network. Due to resource constraints, we cannot validate this prediction on a real cluster of this size.
@@ 7. Conclusion
In this work, we present a novel framework, NeuSight, that forecasts the performance of deep learning execution on GPU servers. In developing NeuSight, we leverage the execution deep learning kernel on a GPU architecture such as tiled execution and use this insight to decompose the prediction problem to smaller sub-problems, i.e., per-tile prediction. For each tile, we use machine learning models to capture the complex non-linear relationship between achieved throughput and input features, and impose performance laws to bound the predicted latency. Results on a variety of models and GPUs, show that NeuSight can provide better predictions than prior works in the area. We open-sourced NeuSight to benefit broader research community.  [^1] NeuSight is available at github.com/sitar-lab/NeuSight
8. Acknowledgements
We thank the anonymous reviewers for their insightful comments. This research was supported through computational resources provided by Partnership for an Advanced Computing Environment (PACE) at Georgia Tech, Institute for Data Engineering and Science (IDEaS) at Georgia Tech, Google Cloud, and AMD HPC Fund Research Cloud. This work is partially supported by Gifts from Google and AMD. The views and conclusions contained herein are those of the authors. They should not be interpreted as representing the official policies or endorsements, either expressed or implied, of Georgia Tech or Meta.
Appendix A Artifact Appendix
A.1. Abstract
The artifact includes source code, Python scripts, and prediction datasets required to reproduce the error rates of different ML models on various GPUs, as described in the Evaluation section (Section 6).
A.2. Artifact check-list (meta-information)
  • Algorithm: ML-based GPU execution time predictor
  • Program: Python, PyTorch
  • Model: Transformer models from HuggingFace
  • Data set: Measured latencies of single kernels and end-to-end ML model executions collected from 8 NVIDIA GPUs and 3 AMD GPUs (available in the GitHub repository)
  • Run-time environment:
    • Ubuntu 20.04, Python 3.9, PyTorch 2.1.0 or similar
    • (If collecting datasets from scratch) PyTorch 2.1 + CUDA 12.1 and PyTorch 2.4 + RoCM 6.1
  • Hardware:
    • Any GPU supporting PyTorch 2.1 or similar
    • (If collecting datasets from scratch) P4, P100, V100, T4, A100, L4, H100, MI100, MI200, and MI250
  • Execution: (If collecting datasets from scratch) Sole GPU usage recommended for profiling
  • Metrics: Predicted latency and prediction error
  • Output: CSV-formatted predicted latency and prediction error values used for tables and figures in the Evaluation section (Section 6)
  • Disk space required: ~50 GB
  • Time to prepare workflow: ~1 hour
  • Time to complete experiments: ~1 hour using the provided dataset; ~10 hours to collect datasets, labels, and traces from scratch
  • Publicly available: Yes
A.3. Description
A.3.1. How to access
The scripts and code are available at https://github.com/sitar-lab/NeuSight. Documentation and setup instructions for artifact evaluation are included in scripts/asplos/README.
A.3.2. Hardware dependencies
Experiments were conducted on P4, P100, V100, T4, A100, L4, H100, MI100, MI200, and MI250. If using the provided dataset, any GPU supporting PyTorch 2.1 or similar is sufficient.
A.3.3. Software dependencies
  • For NVIDIA GPUs: Ubuntu 20.04 or similar, Python 3.9, PyTorch 2.1.0, CUDA Toolkit 12.1
  • For AMD GPUs: RHEL 9.4 or similar, Python 3.12.5, PyTorch 2.4.1, RoCM 6.1
A.3.4. Data sets
The publicly available dataset included in the GitHub repository contains operator latencies and ground truth latency measurements of various ML models across different GPUs.
A.4. Installation and Testing
A.4.1. Installation
git clone https://github.com/sitar-lab/NeuSight
cd NeuSight
pip install -e .
A.4.2. Basic Test
bash scripts/example/gpt3_inference_h100.sh
The printed output should be approximately ~670 ms.
A.5. Experiment workflow
1. (Optional) Run scripts to collect datasets from scratch.
2. (Optional) Train NeuSight with the collected dataset.
3. Use the provided evaluation scripts to predict model execution latencies on GPUs evaluated in the paper.
4. Compare predictions against ground truth using the provided summary.py and table.py scripts.
A.6. Evaluation and expected results
The artifact reproduces the results shown in Figures 7, 8, and 9, and Tables 7 and 8 of the paper. Data for these figures and tables is available in /scripts/asplos/summary. Note that latency predictions may vary slightly (~10%) due to the non-deterministic behavior of DNN models.